OpenStudy (anonymous):

Log8(x+6)-log8(x)=log8(66)

OpenStudy (anonymous):

\[\log_8(x+6)-\log_8x=\log_866\] \[\log_8(\frac{ x+6 }{ x })=\log_866\] using the fact that. \[\log_ab=x\] is the same as writing \[a^x=b\] \[\frac{ x+6 }{ x }=8^{\log_866}\] \[\frac{ x+6 }{ x } = 66\] \[x+6=66x\] \[x=\frac{ 6 }{ 65 }\]

OpenStudy (anonymous):

do you understand what i have done??

OpenStudy (anonymous):

Ya thanks

OpenStudy (anonymous):

yw..

OpenStudy (anonymous):

NICELY SOLVE..............

OpenStudy (anonymous):

thank you...